You are short 15 gasoline futures contracts, established at an initial settle price of $2.66 per gallon, where each contract represents 42,000 gallons. Over the subsequent four trading days, gasoline settles at $2.63, $2.68, $2.71, and $2.76, respectively.

  

Calculate the profit or loss for each trading day. (Do not round intermediate calculations. A negative answer should be indicated by a minus sign.)

      

Profit/Loss
  Day 1 $   
  Day 2 $   
  Day 3 $   
  Day 4 $   

  

Compute your total profit or loss at the end of the trading period. (Do not round intermediate calculations. Input your answer as a positive value.)

  

  (Click to select)ProfitLoss $   
Answer #1

Profit from short forward=number of contracts*contract size*(short price-final price)

1.
=15*42000*(2.66-2.63)=18900.00

2.
=15*42000*(2.63-2.68)=-31500.00

3.
=15*42000*(2.68-2.71)=-18900.00

4.
=15*42000*(2.71-2.76)=-31500.00

5.
=15*42000*(2.66-2.76)=-63000.00